Calculate the price elasticity of demand from point A to B by using the given data below. If the price elasticity of demand│εP│> 1 OR │εP│< 1 (demand curve is elastic/inelastic), discuss why an increase in Price will increase/decrease the total revenue? 

Point Quantity Price

A 500 24

B 400 16


1
Expert's answer
2021-04-29T10:28:20-0400

The price elasticity of demand is calculated as:

"E_d=\dfrac{\Delta Q}{\Delta P}\times \dfrac{P_1+P_2}{Q_1+Q_2}"

If the quantity demanded is 500 when the price is 16 and 400 when the price is 24, then:


"E_d=\dfrac{400-500}{24-16}\times \dfrac{24+16}{500+400}\\[0.4cm]
E_d\approx -0.56"

Because demand is inelastic, the total revenue will increase when the price is increased and decrease when the price is decreased.